How to determine equilibrium level of income. Where


Y=C+Io+Go

C=a+bYd

Yd=Y-T

a=100

Io=100

b=0.75

T=200

Go=200


1
Expert's answer
2021-03-08T09:19:49-0500

Y=C+Io+Go

Y=a+bYd+Io+Go

Y=100+0.75Yd+100+200

Y=0.75Yd+400


Yd=Y-T

Y=Yd+T

Y=Yd+200


Yd+200=0.75Yd+400

Yd-0.75Yd=400-200

0.25Yd=200

Yd=200:0.25

Yd=800


Y=0.75Yd+400=0.75ั…800+400=1000

Y=Yd+200=800+200=1000